QUESTION 2: i) Use (a) no-return payback, (b) discounted payback at 12%, and (c) PW analysis at 12% to select a system. Comment on the results. (10) System 2 System 1 12,000 3,000 8,000 1,000 (year 1-4) 3,000 (year 5-14) 14 First cost, $ NCF, S per year Maximum life, years Solution:
i) What is guideline/Criterion for economic justification of single project using B/C analysis? Is following roject justified? (10) P=$5 AOC $300 K per year $1,2000 K per year B D $200 K per year 1-7% n 10 years
Answer #1
ANALYSIS OF SYSTEM 1
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=12%=0.12
N=Year of Cash Flow
N Year 0 1 2 3 4 5 6 7
CF Cash flow ($12,000) $3,000 $3,000 $3,000 $3,000 $3,000 $3,000 $3,000 SUM
PV=CF/(1.12^N) Present value of Cash Flow ($12,000) $2,679 $2,392 $2,135 $1,907 $1,702 $1,520 $1,357 $1,691
Cumulative Cash Flow(No Return) ($12,000) ($9,000) ($6,000) ($3,000) $0 $3,000 $6,000 $9,000
Cumulative Present Value (12% discounted) ($12,000) ($9,321) ($6,930) ($4,795) ($2,888) ($1,186) $334 $1,691
Payback Period =Period in which Cumulative cash flow=Zero(Period when initial cost is recovered)
(a) No return payback= 4 years
(b) Discounted Payback at 12%=5+(1186/1520)           5.78 years
Net Present worth $1,691
Net Present worth of repeat investment $765 (1691/(1.12^7)
.(c) Total Net present Worth of 14 years operation $2,456 (1691+765)
ANALYSIS OF SYSTEM 2
N Year 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14
CF Cash flow ($8,000) $1,000 $1,000 $1,000 $1,000 $3,000 $3,000 $3,000 $3,000 $3,000 $3,000 $3,000 $3,000 $3,000 $3,000 SUM
PV=CF/(1.12^N) Present value of Cash Flow ($8,000) $893 $797 $712 $636 $1,702 $1,520 $1,357 $1,212 $1,082 $966 $862 $770 $688 $614 $5,810
Cumulative Cash Flow(No Return) ($8,000) ($7,000) ($6,000) ($5,000) ($4,000) ($1,000) $2,000 $5,000 $8,000 $11,000 $14,000 $17,000 $20,000 $23,000 $26,000
Cumulative Present Value (12% discounted) ($8,000) ($7,107) ($6,310) ($5,598) ($4,963) ($3,260) ($1,740) ($383) $828 $1,910 $2,876 $3,738 $4,508 $5,196 $5,810
Payback Period =Period in which Cumulative cash flow=Zero(Period when initial cost is recovered)
(a) No return payback=(5+(1000/30000           5.33 years
(b) Discounted Payback at 12%=7+(383/1212)           7.32 years
Net Present worth $5,810
PW of System 2 is higher . Hence System 2 is recommended
Payback (No return) and discounted payback of System 2 is higher